# RMDM: Radio Map Diffusion Model with Physics Informed

### v2

@@ 3.2.1 UNet 1: Conditional Generation Model with PINN Alignment
The network extracts multi-scale features $\{z_{1},z_{2},\dots,z_{L}\}$ and integrates a physics-informed neural network (PINN) alignment module in the decoder to ensure adherence to physical laws.
@@
$\mathcal{L}_{\text{Cond}}=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}+\lambda_{\text{PINN}}\mathcal{L}_{\text{PINN}}+\lambda_{\text{Reg}}\mathcal{L}_{\text{Reg}},$ (8)
where $\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$ ensures the effectiveness and accuracy of feature extraction by aligning the model output with the real data. $\mathcal{L}_{\text{Reg}}$ represents a regularization term that promotes smoothness or penalizes overfitting, depending on the specific design of the model or the application.
3.2.2 UNet 2: Diffusion Model for Feature Refinement and Map Generation
@@
$\mathcal{L}_{\text{Diff}}=\mathbb{E}_{x_{0},t,\epsilon}[\|\epsilon-\epsilon_{\theta}(x_{t},t,\hat{y})\|^{2}],$ (10)
@@
3.2.3 Overview of Model Architecture
The proposed framework consists of two main components: UNet 1 for feature extraction and UNet 2 for feature refinement and radio map generation. The overall training objective is defined by the total loss function:
$\mathcal{L}_{\text{Total}}=\lambda_{\text{Cond}}\mathcal{L}_{\text{Cond}}+\lambda_{\text{Diff}}\mathcal{L}_{\text{Diff}},$ (11)
@@ 3.2.3 Overview of Model Architecture
$F_{\text{attn}}=\alpha(\text{FFT}(F_{\text{input}}))\cdot F_{\text{spatial}},$ (12)
@@
Algorithm 1 Radio Map Diffusion Model (RMDM)
@@
1:  Initialize U-Net1111($\theta_{1}$), U-Net2222($\theta_{2}$)
@@
4 Experiment
In this study, we evaluate the performance of our proposed model using the RadioMapSeer dataset, which is part of the Pathloss RM Construction Challenge. To validate the effectiveness of our approach, we conducted experiments in three different setups and compared the results with state-of-the-art models, including RME-GAN, RadioUNet, and RadioDiff. The results demonstrate that our model consistently outperforms these methods, achieving superior performance. In addition, we conducted an ablation study to assess the contribution of individual components in our method. As shown in Figure 3, we also provide a visualization comparing our results with those of RadioUNet, highlighting the improved quality and precision of our approach.
4.1 Dataset
The RadioMapSeer dataset Yapar et al. (2022) consists of 700 maps, each representing unique geographic information, such as building data, and includes 80 transmitter locations per map along with corresponding ground truth data. Each map contains between 50 and 150 buildings. We divided the dataset into 500 maps for training and the remaining 200 maps for testing, ensuring no overlapping terrain information between the training and testing sets.
The city maps are sourced from OpenStreetMap and cover metropolitan areas such as Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. In the dataset, the heights of transmitters, receivers, and buildings are set to 1.5 meters, 1.5 meters, and 25 meters, respectively. The transmitter power is configured at 23 dBm, and the carrier frequency is 5.9 GHz. Each map is converted into a 256 × 256 pixel morphological 2D image with a resolution of 1 meter per pixel. Pixels are assigned binary values: ‘1’ for areas inside buildings and ‘0’ for areas outside. Transmitter positions are stored in a two-dimensional numerical format and represented in morphological images, with the transmitter’s pixel set to ‘1’ and all others to ‘0’.
The radio maps in this dataset are generated using software such as WinProp Jakobus et al. (2018). To ensure accurate ground truth for training, the dataset employs Maxwell’s equations to construct radio maps (RMs). Pathloss is calculated by considering the reflection and diffraction of electromagnetic rays. The dataset includes two types of RMs: Static RM (SRM), which considers the impact of static buildings, and Dynamic RM (DRM), which includes both static buildings and randomly generated vehicles along the roads. This comprehensive data set enables a robust evaluation of the performance of our model in realistic urban conditions.
4.2 Experiment Setup
@@
4.3 Results and Analysis
To evaluate the performance of our proposed model, we conducted experiments under three distinct setups: (1) Static Radio Maps (SRM), (2) Dynamic Radio Maps (DRM), and (3) Unbalanced Sample Distribution Among Regions. The results for each setup are detailed below, demonstrating the efficacy of RMDM in comparison to existing state-of-the-art methods.
@@ 4.3 Results and Analysis
RMDM (ours) 0.0031 0.0125 0.978
Table 1: Performance comparison of different models on the SRM setup, reported in terms of NMSE, RMSE, and SSIM. The results highlight the superior performance of RMDM, which achieves the lowest NMSE and RMSE values and the highest SSIM score, demonstrating its capability to generate precise and structurally consistent radio maps. These metrics confirm RMDM’s effectiveness in leveraging physical constraints and diffusion-based modeling to address challenges in static propagation environments.
4.3.1 Setup 1: Static Radio Maps (SRM)
In this configuration, we utilized static building information and transmitter location data as inputs to assess the model’s performance. As presented in Table 1, our proposed model, RMDM, outperforms existing models across all metrics. Specifically, RMDM achieves the lowest NMSE of 0.0031 and RMSE of 0.0125, while attaining the highest SSIM of 0.978. These results indicate that RMDM surpasses state-of-the-art models such as RadioDiff, RadioUNet, and RME-GAN, demonstrating its capability to generate precise and structurally consistent radio maps in static scenarios.
4.3.2 Setup 2: Dynamic Radio Maps (DRM)
To evaluate the model’s performance under dynamic conditions, we incorporated vehicle information into the input data alongside building and transmitter location information. The results in Table 2 illustrate that RMDM consistently outperforms competing models, achieving an NMSE of 0.0047, an RMSE of 0.0146, and an SSIM of 0.968. These metrics demonstrate the robustness of RMDM in modeling dynamic urban environments, where additional complexity is introduced by moving objects such as vehicles.
@@ 4.3.2 Setup 2: Dynamic Radio Maps (DRM)
RMDM (ours) 0.0047 0.0146 0.968
Table 2: Performance comparison of different models on the DRM setup, reported in terms of NMSE, RMSE, and SSIM. The results underscore the robustness of RMDM in dynamic environments, achieving the lowest NMSE and RMSE values and a high SSIM score, showcasing its ability to handle additional complexity introduced by moving objects and varying propagation conditions.
4.3.3 Setup 3: Unbalanced Sample Distribution
In the third setup, we tested the model’s ability to handle unbalanced sample distributions, with sampling ratios varying randomly between 1% and 10%. Table 3 shows that RMDM achieves the best performance among all methods, with an NMSE of 0.0022 and an RMSE of 0.0117. RMDM significantly outperforms traditional interpolation-based methods, such as Kriging and RBF, as well as deep learning approaches like AE, Deep AE, and Unet. These results highlight RMDM’s capability to generate high-quality radio maps even in the presence of sparse and unevenly distributed observations.
4.3.4 Overall Performance
Across all experimental setups, RMDM demonstrates superior performance compared to state-of-the-art methods. Its ability to achieve lower NMSE and RMSE values while maintaining high SSIM scores underscores the effectiveness of our approach in constructing accurate and reliable radio maps under diverse conditions. The results validate the robustness, generalizability, and efficiency of the proposed model, making it a promising solution for radio map construction challenges.
4.4 Ablation Study
@@
RMDM (ours) 0.0022 0.0117
Table 3: Performance comparison of different models for Setup 3, reported in terms of NMSE and RMSE. The results demonstrate the superior performance of RMDM in handling unbalanced sample distributions, achieving the lowest NMSE and RMSE values. This highlights its capability to generate accurate radio maps even under sparse and unevenly distributed observations.
$\mathcal{L}_{\text{MSE}}$ $\mathcal{L}_{\text{PINN}}$ $\mathcal{L}_{\text{Reg}}$ NMSE $\downarrow$ RMSM $\downarrow$
@@
Table 4: Ablation study results showing the impact of $L_{\text{MSE}}$, $L_{\text{PINN}}$, and $L_{\text{Reg}}$ on NMSE and RMSE. The results highlight the critical role of $L_{\text{MSE}}$ in minimizing prediction errors, $L_{\text{PINN}}$ in enforcing physical consistency, and $L_{\text{Reg}}$ in ensuring stability and generalization. The combination of all three loss components achieves the best performance, demonstrating their synergistic effect in enhancing model accuracy and robustness.
These results validate the effectiveness of the proposed loss formulation, demonstrating that integrating all components leads to superior performance in terms of both accuracy and stability. The findings also emphasize that while individual components contribute positively, their combined use ensures a comprehensive balance between error reduction, physical consistency, and regularization, which is critical for accurate and reliable radio map reconstruction.
4.5 Reflection on Evaluation Metrics
In our experiments, we identified certain limitations associated with the commonly used evaluation metrics, such as NMSE and RMSE. These metrics are fundamentally based on the calculation of mean squared error, which averages the prediction errors across the entire dataset. While this approach provides a general measure of accuracy, it may fail to capture critical details in specific scenarios. For instance, when a model produces blurred boundaries in regions of signal intensity, it can still achieve a low MSE, leading to seemingly favorable results Zhang and Wang (2022) . However, such outcomes are evidently misleading, as they fail to account for the detailed variations in the signal distribution, which are essential for accurate radio map reconstruction. Therefore, it is necessary to adopt more fine-grained evaluation methods that can comprehensively assess the model’s ability to preserve spatial details and recover critical boundaries in the signal map Shrestha et al. (2024).
@@ 4.5 Reflection on Evaluation Metrics
Figure 3: Visualization of the results generated by RMDM compared to RadioUnet. The proposed RMDM model demonstrates more fine-grained results and exhibits closer alignment with the ground truth data, highlighting its superiority in capturing intricate details.
5 Conclusion
We present the Radio Map Diffusion Model (RMDM), a physics-informed framework that integrates the Helmholtz equation with a dual U-Net architecture to tackle the challenges of radio map reconstruction in complex propagation environments. By embedding Physics-Informed Neural Networks (PINNs) into the learning process, we enforce physical consistency and leverage diffusion-based refinement to generate highly accurate and robust radio maps, even with sparse and uneven data. Our extensive experiments demonstrate that RMDM consistently outperforms state-of-the-art methods across static, dynamic, and unbalanced scenarios, achieving significant improvements in NMSE, RMSE, and SSIM. These results highlight our model’s ability to effectively combine physics-based constraints with data-driven learning, setting a new paradigm for radio map reconstruction. While our approach demonstrates superior performance, the dual U-Net architecture introduces computational complexity, and hyperparameter sensitivity remains a challenge. In the future, we aim to improve computational efficiency, adapt the model to dynamic propagation conditions, and extend its applicability to real-time and large-scale scenarios. Furthermore, we plan to integrate multi-source data and enhance interpretability through advanced visualization techniques, paving the way for practical applications in 5G/6G network optimization, intelligent transportation systems, and spectrum management.
Acknowledgements
This work was supported by Guangzhou-HKUST(GZ) Joint Funding ProgramGrant No.2023A03J0008, Education Bureau of Guangzhou Municipality.

### v3

@@ 3.2.1. UNet 1: Conditional Generation Model with PINN Alignment
The network extracts multi-scale features $\{z_{1},z_{2},\dots,z_{L}\}$ and incorporates a representation alignment module in the decoder, which ensure adherence to physical laws while prioritizing alignment with pre-trained radio signal representations.
@@
(8) $\mathcal{L}_{\text{Cond}}=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}+\lambda_{\text{PINN}}\mathcal{L}_{\text{PINN}}+\lambda_{\text{Reg}}\mathcal{L}_{\text{Reg}},$
where $\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i}(\hat{y}_{i}-y_{i})^{2}$ensures the effectiveness and accuracy of feature extraction by aligning the model output with the real data. $\mathcal{L}_{\text{Reg}}$ represents a regularization term that promotes smoothness or penalizes overfitting, depending on the specific design of the model or the application.
3.2.2. UNet 2: Diffusion Model for Feature Refinement and Map Generation
@@
(10) $\mathcal{L}_{\text{Diff}}=\mathbb{E}_{x_{0},t,\epsilon}[\|\epsilon-\epsilon_{\theta}(x_{t},t,\hat{y})\|^{2}],$
@@
3.2.3. Overview of Model Architecture
The proposed framework consists of two main components: UNet 1 for feature extraction and representation alignment and UNet 2 for feature refinement and radio map generation. The overall training objective is defined by the total loss function:
(11) $\mathcal{L}_{\text{Total}}=\lambda_{\text{Cond}}\mathcal{L}_{\text{Cond}}+\lambda_{\text{Diff}}\mathcal{L}_{\text{Diff}},$
@@ 3.2.3. Overview of Model Architecture
(12) $F_{\text{attn}}=\alpha(\text{FFT}(F_{\text{input}}))\cdot F_{\text{spatial}},$
@@
Algorithm 1 Physics-Aligned Radio Map Diffusion Model (Phy-RMDM)
@@
1: Initialize U-Net11($\theta_{1}$), U-Net22($\theta_{2}$)
@@
4. Experiment
This study evaluates PhyRMDM, a framework designed to overcome the inherent challenges of applying physical constraints to diffusion-based generative models for radio map reconstruction. Using the RadioMapSeer dataset from the Pathloss RM Construction Challenge, we validate our approach across three distinct experimental setups, comparing its performance against state-of-the-art models like RME-GAN, RadioUNet, and RadioDiff. The results demonstrate that our model consistently outperforms these methods. We also conduct a detailed ablation study to assess the contribution of individual components. As shown in Figure 4, we provide a visual comparison with RadioUNet, highlighting the improved quality and precision of our approach.
4.1. Dataset
The RadioMapSeer dataset (Yapar et al., 2022) contains 700 maps with unique geographic data, including 50–150 buildings and 80 transmitter locations per map, plus ground truth. We use 500 maps for training and 200 for testing with no terrain overlap.
Maps from OpenStreetMap cover cities such as Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. Transmitter, receiver, and building heights are set to 1.5 m, 1.5 m, and 25 m, respectively, with transmitter power at 23 dBm and carrier frequency at 5.9 GHz. Each map is a 256×256 binary image (1 m/pixel): ‘1’ for building interiors, ‘0’ otherwise; transmitter positions are similarly encoded with the transmitter pixel set to ‘1’.
The radio maps in this dataset are generated using software such as WinProp (Jakobus et al., 2018). To ensure accurate ground truth for training, the dataset employs Maxwell’s equations to construct radio maps (RMs). Pathloss is calculated by considering the reflection and diffraction of electromagnetic rays. The dataset includes two types of RMs: Static RM (SRM), which considers the impact of static buildings, and Dynamic RM (DRM), which includes both static buildings and randomly generated vehicles along the roads. This comprehensive data set enables a robust evaluation of the performance of our model in realistic urban conditions.
4.2. Experiment Setup
@@
Figure 3. Visual comparison of ablation study on sparse radio map reconstruction. Each triplet shows results from PhyRMDM with PINN, without PINN, and ground truth. PINN enhances structural details and accuracy, highlighting its effectiveness in physics-guided reconstruction.
4.3. Results and Analysis
To evaluate the performance of our proposed model, we conducted experiments under three distinct setups: (1) Static Radio Maps (SRM), (2) Dynamic Radio Maps (DRM), and (3) Unbalanced Sample Distribution Among Regions. The results for each setup are detailed below, demonstrating the efficacy of PhyRMDM in comparison to existing state-of-the-art methods.
@@ 4.3. Results and Analysis
PhyRMDM (ours) 0.0031 0.0125 0.978
Table 1. Performance comparison of different models on the SRM setup in terms of NMSE, RMSE, and SSIM. PhyRMDM achieves the lowest NMSE/RMSE and highest SSIM, demonstrating its ability to generate accurate and structurally consistent radio maps. These results confirm its effectiveness in combining physical constraints with diffusion-based modeling for static propagation environments.
4.3.1. Setup 1: Static Radio Maps (SRM)
In this configuration, we utilized static building information and transmitter location data as inputs to assess the model’s performance. As presented in Table 1, our proposed model, PhyRMDM, outperforms existing models across all metrics. Specifically, PhyRMDM achieves the lowest NMSE of 0.0031 and RMSE of 0.0125, while attaining the highest SSIM of 0.978. These results indicate that PhyRMDM surpasses state-of-the-art models such as RadioDiff, RadioUNet, and RME-GAN, demonstrating its capability to generate precise and structurally consistent radio maps in static scenarios.
4.3.2. Setup 2: Dynamic Radio Maps (DRM)
To evaluate the model’s performance under dynamic conditions, we incorporated vehicle information into the input data alongside building and transmitter location information. The results in Table 2 illustrate that PhyRMDM consistently outperforms competing models, achieving an NMSE of 0.0047, an RMSE of 0.0146, and an SSIM of 0.968. These metrics demonstrate the robustness of PhyRMDM in modeling dynamic urban environments, where additional complexity is introduced by moving objects such as vehicles.
@@ 4.3.2. Setup 2: Dynamic Radio Maps (DRM)
PhyRMDM (ours) 0.0047 0.0146 0.968
Table 2. Performance comparison of different models on the DRM setup in terms of NMSE, RMSE, and SSIM. PhyRMDM shows strong robustness in dynamic environments, achieving the lowest NMSE/RMSE and high SSIM, highlighting its ability to manage complexity from moving objects and varying propagation conditions.
4.3.3. Setup 3: Unbalanced Sample Distribution
In the third setup, we tested the model’s ability to handle unbalanced sample distributions, with sampling ratios varying randomly between 1% and 10%. Table 3 shows that PhyRMDM achieves the best performance among all methods, with an NMSE of 0.0022 and an RMSE of 0.0117. PhyRMDM significantly outperforms traditional interpolation-based methods, such as Kriging and RBF, as well as deep learning approaches like AE, Deep AE, and Unet. These results highlight PhyRMDM’s capability to generate high-quality radio maps even in the presence of sparse and unevenly distributed observations.
4.3.4. Overall Performance
Across all setups, PhyRMDM outperforms state-of-the-art methods, achieving lower NMSE and RMSE with higher SSIM. These results highlight its robustness, generalizability, and efficiency, demonstrating its effectiveness for accurate and reliable radio map construction.
4.4. Ablation Study
@@
PhyRMDM (ours) 0.0022 0.0117
Table 3. Performance comparison of different models for Setup 3 using NMSE and RMSE. PhyRMDM achieves the lowest errors, showing its strength in handling unbalanced sample distributions and generating accurate radio maps under sparse, uneven observations.
$\mathcal{L}_{\text{MSE}}$ $\mathcal{L}_{\text{PINN}}$ $\mathcal{L}_{\text{Reg}}$ NMSE $\downarrow$ RMSE $\downarrow$
@@
Table 4. Ablation results showing the effects of \$L_MSE\$, \$L_PINN\$, and \$L_Reg\$ on NMSE and RMSE. \$L_MSE\$ reduces prediction errors, \$L_PINN\$ enforces physical consistency, and \$L_Reg\$ improves stability and generalization. Combining all three yields the best performance, demonstrating their synergistic effect on accuracy and robustness.
Figure 3 compares results with and without $\mathcal{L}\text{PINN}$. Without it, models produce blurred boundaries and oversmoothed transitions near buildings, failing to capture diffraction and reflection due to missing Helmholtz-based guidance. Adding $\mathcal{L}\text{PINN}$ enhances structural fidelity and enables fine-grained, physically consistent predictions even with sparse data, highlighting its importance for realistic signal modeling.
Overall, combining all loss components yields the best accuracy and stability, balancing error reduction, physical consistency, and regularization for reliable radio map reconstruction.
@@
Figure 4. Visualization of the results generated by PhyRMDM compared to RadioUnet. The proposed PhyRMDM model demonstrates more fine-grained results and exhibits closer alignment with the ground truth data, highlighting its superiority in capturing intricate details.
5. Conclusion
We introduce the Physics-Informed Radio Map Diffusion Model (PhyRMDM), a framework for radio map reconstruction that integrates physics-based constraints with generative models. Directly constraining diffusion models is difficult due to complex radio environments and noise prediction. PhyRMDM tackles this with a dual U-Net design: one U-Net enforces physical alignment using an approximate Helmholtz equation, while the other refines the output through a parallel diffusion process. This approach produces accurate, robust radio maps even with sparse data. Experiments show PhyRMDM outperforms state-of-the-art methods across static, dynamic, and unbalanced scenarios, with notable improvements in NMSE, RMSE, and SSIM. While effective, the dual-U-Net structure is computationally heavy and sensitive to hyperparameters. Future work will focus on improving efficiency, adapting to dynamic conditions, and enabling real-time, large-scale applications with multi-source data and better interpretability for 5G/6G network optimization and spectrum management.